What is the value of tan15. 

Grade:12

5 Answers

Chetan Mandayam Nayakar
10 years ago

let tan15=x,tan30=sqrt(1/3),

2x/(1-x2)=tan30

Shivam Dimri
10 years ago

use the formula tan A-B

a=45 and b= 30

i prefer if you solve urself..

Sameer Gupta
10 years ago

Dear preet,

               tan 15=tan(45-30)

                        =(tan 45-tan 30)/(1+tan 45*tan 30)

                          [As tan(A-B)=(tan A-tan B)/(1+tan A*tan B)}

                        giving value : tan 45=1 and tan 30=1/underroot 3

                        we get,tan 15=({underroot 3}-1)/({underroot 3}+1)

                                           =1.732-1/1.732+1

                                           =0.732/2.732

                                           =0.26794

                                           =0.268.------>> ANS:

                                            

harihara kumar
10 years ago

0.2679

Akash Kumar Dutta
10 years ago

tan(45-30)= root3 - 1 / root3 + 1= [root3-1]^2 / 2 = 2-root3 ANS.